OpenStudy (harlequin):

Can anyone show me how to work these out? Solve each equation in the complex number system. X^4=1 X^4+3x^2-4=0

hartnn (hartnn):

i can get you started : 1. x^4 = 1 (x^2)^2 -1^2 = 0 (x^2+1)(x^2-1) =0 .... 2. x^4+3x^2 -4=0 Let x^2 = a a^2 +3a-4 = 0 find 2 values of a first. Lets see how much you can work each of them out. Go!

OpenStudy (harlequin):

Thank you so much that actually helped. I always have trouble getting to that point.

hartnn (hartnn):

welcome ^_^